OpenStudy (anonymous):

A ball at rest starts rolling down a hill with a constant acceleration of 3.2 meters/second2. What is the final velocity of the ball after 6.0 seconds?

OpenStudy (abhisar):

Here, Initial velocity (u)=0 m/s Acceleration (a) = 3.2 m/\(\sf s^2\) Time = 6 seconds We can use the equation v=u+at to find the final velocity (v)